Assume you travel a lot as part of your business and therefore you drive 60,000 miles each year. your car manual recommends oil every 7,500 miles but also recommends a more expensive oil priced at $50 per oil change. your service center recommends cheaper oil priced at $20 per oil change but also recommends changing oil at every 5,000 miles.Assuming price of labor is $15 per oil change and ignoring all other costs.how much money will you save each year if you follow the recommendation of your service center,instead of your car manual.
Answer #1

Answer:

Annual cost of changing oil:

(i) As per recommendation of car manual:

Total miles driven per year = 60,000

Car manual recommends oil every 7,500 miles priced at $50 per oil change

Price of labor is $15 per oil change

Number of times 'change of oil' required per year = 60000 / 7500 = 8

Annual cost of changing oil = 8 * ($50 + $15) = $520

(ii) As per recommendation of your service center:

Service center recommends oil every 5000 miles priced at $20 per oil change

Price of labor is $15 per oil change

Number of times 'change of oil' required per year = 60000 / 5000 = 12

Annual cost of changing oil = 12 * ($20 + $15) = $420

Hence:

Money you will save each year if you follow the recommendation of your service center, instead of your car manual = 520 - 420 = $100

Money you will save each year if you follow the recommendation of service center, instead of your car manual = $100